OpenStudy (anonymous):

It rains in Spain once every 10 days, and when it does there is a 2% chance of hurricane in Hartford. When it does not rain in Spain, there is a 1% chance of hurricane in Hartford. What is the probability that it rains in spain when there is a hurricane in Hartford?

OpenStudy (ybarrap):

Let RS = It rains in Spain (and NRS means no rain in Spain) Let HH = There is a hurricane in Hartford We know $$ P(RS)=\cfrac{1}{10}\\ P(NRS)=1-P(RS)=\cfrac{9}{10}\\ P(HH|RS)=\cfrac{2}{100}\\ P(HH|NRS)=\cfrac{1}{100}\\ P(HH)=P(HH|RS)P(RS)+P(HH|NRS)P(NRS)\\ =\cfrac{2}{100}\cfrac{1}{10}+\cfrac{1}{100}\cfrac{9}{10} $$ What we need is the probability of RS given HH: $$ P(RS|HH)=\cfrac{P(HH|RS)P(RS)}{P(HH)} $$ And we know all three probabilities on the right. Does this make sense?